OpenStudy (anonymous):

solve. 2(6 - 2n) -3(n - 1) = 5(n - 1)

OpenStudy (anonymous):

n=2

OpenStudy (anonymous):

The answer is 5/3. I was just wondering why.

OpenStudy (anonymous):

12-4n-3n-3=5n-5 12-4n-3=5n-5 14=7n n=2

OpenStudy (anonymous):

thats the works

OpenStudy (anonymous):

It helped me. Thanks.